# ASURA-FDPS-ML: Star-by-star Galaxy Simulations Accelerated by Surrogate Modeling for Supernova Feedback

### v1

@@ 5.2 Phase Structures of the Wind
Literatures on the simulations of even heavier galaxies also show the transition between sub-sonic wind at low altitude ($z\sim 1$ kpc) and super-sonic wind at high altitude ($z>10$ kpc) in 2D PDFs likely caused by the fountain. Steinwandel et al. (2024a, b), which used an isolated galaxy with ten times heavier virial mass than this paper, showed the sub-to-super sonic transition of hot winds between 1 kpc and 10 kpc. Pandya et al. (2021) using one hundred times heavier galaxy than this paper, showed the hot super-sonic wind at $\sim 25$ kpc. The majority of the hot wind at the high altitude is supersonic, while the subsonic gas drops off.
@@ Acknowledgments
The authors thank Carolina Cuesta-Lazaro, Chang-Goo Kim, Eve Ostriker, and Philip F. Hopkins for fruitful discussions. This work was supported by JSPS KAKENHI Grant Numbers 22H01259, 22KJ0157, 21H04499, 21K03614, 23K03446, and 24K07095, and MEXT as “Program for Promoting Research on the Supercomputer Fugaku” (Structure and Evolution of the Universe Unraveled by Fusion of Simulation and AI; Grant Number JPMXP1020230406; Project ID: hp230204, hp240219). KH is supported by JSPS Research Fellowship for Young Scientists and accompanying Grants-in-Aid for JSPS Fellows (22J23077 and 22KJ1153), JEES $\cdot$ Mistubishi corporation science technology student scholarship in 2022, and The University of Tokyo doctoral fellowship with the IIW program, and thanks CCA at the Flatiron Institute for hospitality and the National Science Foundation under Cooperative Agreement PHY-2019786 (The NSF AI Institute for Artificial Intelligence and Fundamental Interactions, http://iaifi.org/) for support while a portion of this research was carried out. YH is supported by the grant OISE-1927130: The International Research Network for Nuclear Astrophysics (IReNA), awarded by the US National Science Foundation (NSF). UPS is supported by a Flatiron Research Fellowship (FRF) at the Flatiron Institutes Center for Computational Astrophysics (CCA). The Flatiron Institute is supported by the Simons Foundation.
@@ Appendix A Convergence Test of our SPH code
Appendix B Surrogate Modeling
We adopt U-Net, a model based on convolutional neural networks (CNNs). We first map SPH particles into voxels (uniform mesh in 3D) with interpolation, which is necessary to process the data with CNNs. After predicting the results, we sampled SPH particle data from the voxels. Such methods have been applied to hydrodynamical simulations (Hirashima et al., 2023a, b; Chan et al., 2024) and $N$-body simulations (He et al., 2019; Jamieson et al., 2023, 2024; Legin et al., 2024). In this section, we describe the interpolation method and our ML model in detail.
B.1 Voxel Generation
@@
$f_{i}=\sum_{j}m_{j}\frac{f_{j}}{\rho_{j}}W(r_{ij},h),$ (B1)
where $m_{j}$, $\rho_{j}$, $r_{ij}$, $h$, and $W(r_{ij},h)$ are the mass, density, distance between the particles $i$ and $j$, and a smooth function. The number of neighboring particles is set to 64. Practically, we also applied a normalized interpolation based on Sheperd’s method (inverse-distance-weighted interpolation; Shepard, 1968; Price, 2007) to improve the precision of quantities interpolated into voxels by mitigating effects related to particle distribution. The normalization is achieved by ensuring that the sum of the interpolation weights $w(i)$ is equal to unity. The weight $w(i)$ of particle $i$ for Equation (B1) is defined as the following:
@@
$=\sum_{j=1}^{N}\frac{m_{j}}{\rho_{j}}W_{i,j}.$ (B2)
The normalized quantities $\tilde{f}(i)$ (i.e., normalized temperature and velocities) can be obtained by dividing Equation (B1) with Equation (B2) as
$\tilde{f}(i):=f(i)/w(i).$ (B3)
@@
$v_{\cdot,p}^{*}$ :=$:=\begin{cases}\log_{10}\tilde{v}_{\cdot}&\text{if }\tilde{v}_{\cdot}>0\\ 0&\text{if }\tilde{v}_{\cdot}\leq 0\end{cases}$ (B6)
$v_{\cdot,n}^{*}$ :=$:=\begin{cases}0&\text{if }\tilde{v}_{\cdot}\geq 0\\ \log_{10}(-\tilde{v}_{\cdot})&\text{if }\tilde{v}_{\cdot}<0\end{cases}$ (B7)
@@
$\hat{y}=\mathcal{M}(X\mid\theta).$ (B8)
B.3 Sampling

### v2

@@ 5.2 Phase Structures of the Wind
Literature on the simulations of even heavier galaxies also show the transition between sub-sonic wind at low altitude ($z\sim 1$ kpc) and super-sonic wind at high altitude ($z>10$ kpc) in 2D PDFs likely caused by the fountain. Steinwandel et al. (2024a, b), which used an isolated galaxy with ten times heavier virial mass than this paper, showed the sub-to-super sonic transition of hot winds between 1 kpc and 10 kpc. Pandya et al. (2021) using one hundred times heavier galaxy than this paper, showed the hot super-sonic wind at $\sim 25$ kpc. The majority of the hot wind at the high altitude is supersonic, while the subsonic gas drops off.
@@ 6 Conclusion and Outlook
While we have focused on the SNe remnants after Sedov phase, the problem of the short timesteps also appears in the earlier phase of free expansion. Such expansion is terminated when the swept-up mass becomes equal to the ejecta mass of the remnant of a few solar masses (e.g., Reynolds et al., 2008), and the reverse shock decouples from the forward shock, propagating backward into the vacuum generated by the forward shock. As the reverse shock heats the gas inside the cavity generated by the forward shock to temperatures of $\sim 10^{7}$ K, the injection of the SN-energy has been generally handled as a single thermal injection event so far as a “subgrid model.” To resolve this phase more accurately, one would need to have at least 10 to 100 times higher mass resolution than the ejecta mass. The kinetic injection schemes have been applied in many successful simulation endeavors of the ISM (e.g., Hu et al., 2016), but this scheme does also require not only a high mass resolution but also a sufficient temporal resolution. The machine learning method similar to the one developed in this study may also prove effective in modeling these phases.
@@ Acknowledgments
The authors thank Carolina Cuesta-Lazaro, Chang-Goo Kim, Eve Ostriker, and Philip F. Hopkins for fruitful discussions. This work was supported by JSPS KAKENHI Grant Numbers 22H01259, 22KJ0157, 25K01046, 21K03614, 22J23077, 22KJ1153, 23K03446, and 24K07095, 23K20035, 24H00004, and MEXT as “Program for Promoting Research on the Supercomputer Fugaku” (Structure and Evolution of the Universe Unraveled by Fusion of Simulation and AI; Grant Number JPMXP1020230406; Project ID: hp230204, hp240219). KH is supported by JSPS Research Fellowship for Young Scientists, JEES $\cdot$ Mitsubishi corporation science technology student scholarship in 2022, The University of Tokyo doctoral fellowship with the IIW program, and the JSPS International Leading Research (ILR) program (JP22K21349), and thanks CCA at the Flatiron Institute for hospitality and the National Science Foundation under Cooperative Agreement PHY-2019786 (The NSF AI Institute for Artificial Intelligence and Fundamental Interactions, http://iaifi.org/) for support while a portion of this research was carried out. YH is supported by the grant OISE-1927130: The International Research Network for Nuclear Astrophysics (IReNA), awarded by the US National Science Foundation (NSF). UPS is supported by a Flatiron Research Fellowship (FRF) at the Flatiron Institutes Center for Computational Astrophysics (CCA). The Flatiron Institute is supported by the Simons Foundation.
@@ Appendix A Convergence Test of our SPH code
Appendix B Pre-evolved Initial Condition
Figure 12: Phase structures of SN-DT (left) as listed in Table 1 at $t=0$, $t=10$ Myr, and $t=100$ Myr. Density-temperature phase diagrams of mass.
While initial condition generators for disk-galaxy simulations (e.g., Springel, 2005) can reconstruct summary statistics, they generally create smoothed gas distribution without detailed structures such as filaments with strong density contrasts and large bubbles created by SNe. The smoothed distribution could cause unrealistic starbursts. To avoid this issue, simulations typically require hundreds of megayears to develop turbulent fields and density contrasts driven by stellar feedback, which means early results must be discarded – a significant computational waste. In this study, to save computational costs, we instead use a snapshot from Steinwandel & Goldberg (2025) that was generated using different codes and cooling functions.
During 500 Myr of pre-evolution, only a small amount of gas mass is converted into stellar mass; the resulting stellar and gas mass at 500 Myr are $2\times 10^{7}~{}\mathrm{M_{\odot}}$ and $3.96\times 10^{7}~{}\mathrm{M_{\odot}}$, respectively. The gas mass change is only 0.67 percent, which is negligible for evolutionary comparisons. While preserving the overall mass distribution of the initial galaxy disk, the pre-simulated galaxy exhibits large bubbles and small-scale density contrasts in the ISM that help prevent artificial starbursts.
For consistency, our simulations do not inherit stars born during the pre-simulation; all SNe in our simulations originate from stars formed within our simulation timeframe. This treatment potentially impacts the ISM phase structure during the transition between numerical solvers. Figure 12 compares ISM phase structures at $t=0$, 10 Myr, and 100 Myr, where the $t=0$ snapshot corresponds to the $t=500$ Myr snapshot in Steinwandel & Goldberg (2025). Between $t=0$ and $t=10$ Myr, we find a reduction in gas with a temperature of $T<10^{2}$ K and a number density of $n>1~{}\mathrm{cm^{-3}}$. The gas kept missing even at $t=100$ Myr, which is likely due to changes in cooling functions. The dense gas undergoes strong cooling in the early period of our simulation, resulting in an artificial starburst as shown in Figure 4. At $t=10$ Myr, fewer gas particles exhibit temperatures above $10^{5}$ K compared to $t=0$, since we did not incorporate stellar feedback from massive stars formed during pre-simulation. These effects, however, diminish by $t=100$ Myr, after the artificial initial starburst subsides and SN feedback effects become prominent.
We conclude that the ISM phase structure is initially affected by the sudden change in the thermodynamic model and the temporary absence of stellar feedback. However, this deficiency in high-temperature gas is largely mitigated following the artificial initial starburst period, when SN feedback effects become prominent.
Appendix C Surrogate Modeling
We adopt U-Net, a model based on convolutional neural networks (CNNs). We first map SPH particles into voxels (uniform mesh in 3D) with interpolation, which is necessary to process the data with CNNs. After predicting the results, we sampled SPH particle data from the voxels. Such methods have been applied to hydrodynamical simulations (Hirashima et al., 2023a, b; Chan et al., 2024) and $N$-body simulations (He et al., 2019; Jamieson et al., 2023, 2025; Legin et al., 2024). In this section, we describe the interpolation method and our ML model in detail.
C.1 Voxel Generation
@@
$f_{i}=\sum_{j}m_{j}\frac{f_{j}}{\rho_{j}}W(r_{ij},h),$ (C1)
where $m_{j}$, $\rho_{j}$, $r_{ij}$, $h$, and $W(r_{ij},h)$ are the mass, density, distance between the particles $i$ and $j$, and a smooth function. The number of neighboring particles is set to 64. Practically, we also applied a normalized interpolation based on Sheperd’s method (inverse-distance-weighted interpolation; Shepard, 1968; Price, 2007) to improve the precision of quantities interpolated into voxels by mitigating effects related to particle distribution. The normalization is achieved by ensuring that the sum of the interpolation weights $w(i)$ is equal to unity. The weight $w(i)$ of particle $i$ for Equation (C1) is defined as the following:
@@
$=\sum_{j=1}^{N}\frac{m_{j}}{\rho_{j}}W_{i,j}.$ (C2)
The normalized quantities $\tilde{f}(i)$ (i.e., normalized temperature and velocities) can be obtained by dividing Equation (C1) with Equation (C2) as
$\tilde{f}(i):=f(i)/w(i).$ (C3)
@@
$v_{\cdot,p}^{*}$ :=$:=\begin{cases}\log_{10}\tilde{v}_{\cdot}&\text{if }\tilde{v}_{\cdot}>0\\ 0&\text{if }\tilde{v}_{\cdot}\leq 0\end{cases}$ (C6)
$v_{\cdot,n}^{*}$ :=$:=\begin{cases}0&\text{if }\tilde{v}_{\cdot}\geq 0\\ \log_{10}(-\tilde{v}_{\cdot})&\text{if }\tilde{v}_{\cdot}<0\end{cases}$ (C7)
@@
$\hat{y}=\mathcal{M}(X\mid\theta).$ (C8)
C.3 Sampling